OpenStudy (anonymous):

solve the system by the substitution method 1/2x+1/5y=-11/10 1/2x+2y=5/2

OpenStudy (shayaan_mustafa):

is it difficult?

OpenStudy (anonymous):

Am trying to compare what i got if you can help!

OpenStudy (anonymous):

Thanx

OpenStudy (shayaan_mustafa):

sure. i can.

OpenStudy (radar):

I would proceed by multiplying the second equation by 2. Getting: x+4y=5 Then express x in terms of y like: x=-4y+5. Substitute for x in the first equation: (1/2)(-4y+5) +1/5y = -11/10 -2y+5/2 +1/5y=-11/10 (-9/5)y=-11/10-5/2 (-9/5)y=-36/10 y=2

OpenStudy (radar):

x+8=5 x=-3 (-3,2)

OpenStudy (anonymous):

Thanx

OpenStudy (radar):

How did this compare? with your answer

OpenStudy (radar):

You're most welcome.

OpenStudy (anonymous):

I had the same (-3, 2) I wasn't sure

OpenStudy (radar):

O.K. probably is correct.